### v2

# BondMatcher: H-Bond Stability Analysis in Molecular Systems

Abstract
This is an abstract, line1.
line 2.
line 3.
line 4.
line 5.
line 6.
line 7.
line 8.
line 9.
line 10.
line 11.
line 12.
Abstract
[…]
keywords:
Quantum chemistry, topological data analysis, discrete Morse theory, ensemble data.
[…]
H-bond stability analysis under vibrational displacements for $4$ isomers of the water hexamer: (a) the Ring (vibration mode 00), (b) the Book (vibration mode 00), (c) the Cage (vibration mode $18$), and (d) the Prism (vibration mode 00). The $6$ water molecules are shown with transparent isosurfaces. Top: Electron density extremum graphs for the input displacements (left, black to red), partial isomorphisms to the equilibrium (right, colored arcs, white: unmatched arcs). Bottom: Equilibrium extremum graphs, colored by bond occurrence rate (black to white to red). Our analysis identifies H-bonds paths (arrows) susceptible of breaking under molecular vibrations (e.g., turning the Cage into the Book, white bond). Our per-bond matching enables a detailed analysis of bonding indicators through the ensemble, corroborating experimental observations and refining geometrical criteria for the disappearance of H-bonds.
Introduction
In chemistry, the properties of a molecular system are determined by the interactions between the atoms forming the system. Two main types of interactions are typically distinguished. First, covalent bonds, refer to interactions that arise from the sharing of electrons between atoms [89]. Second, non-covalent interactions denote all other types of chemical interactions. These interactions are more subtle but they play a crucial role in many biological processes and chemical design tasks (e.g., molecular folding, protein docking). Among the documented non-covalent bonds [76], hydrogen bonds (or H-bonds) [1] form a prominent bond type which intervenes in many molecular systems, from simple water dimers to macromolecules such as DNA.
Nevertheless, H-bonds are challenging to study as they are weaker than covalent bonds. As a consequence, they can easily rearrange under molecular motions such as vibrations or rotations. Then, the condition of appearance and disappearance of H-bonds under such structural rearrangements is an active research topic, both theoretically and experimentally [24]. Such a flexibility challenges the structural description of H-bonds, which often relies on simple, distance-based, heuristic criteria for establishing their presence, without explicit consideration of the underlying electronic effects [1].
In contrast, the Quantum Theory of Atoms In Molecules (QTAIM) [3, 50] provides a framework to describe molecular interactions at the electronic level thanks to the analysis of the electron density field. QTAIM, combined with robust computational tools based on discrete Morse theory (DMT) [23, 5, 79, 7], enables a quantitative, reliable and interpretable analysis of chemical interactions [37, 57]. Still, to the best of our knowledge, these robust computational tools based on DMT have only been used for the static analysis of molecular systems, in particular at their equilibrium state. However, considering the natural dynamic behavior of molecular systems (e.g., vibrations, rotations) is of paramount importance, in particular to assess the stability of H-bonds.
In this work, we fill this gap and extend the topological techniques based on DMT and QTAIM [5, 79, 7] to characterize the dynamic behavior of molecular systems. For this, we introduce a new framework, called BondMatcher, which robustly and automatically matches bonds across multiple dynamic states of a given molecular system. At a technical level, this approach exploits the specific hypotheses about the molecular system to reliably compute relevant partial isomorphisms between the extremum graphs of the electron density fields (Sec. 4). This automatic correspondence enables a fine, quantitative and interpretable analysis of the presence of H-bond paths within molecular systems under various dynamical effects. Specifically, we document the analysis of two case studies focusing on four prominent, low-energy isomers (namely, the Ring, the Book, the Cage and the Prism) of water hexamer, a three-dimensional system playing a central role for the understanding of the different states of water. The first case study (Sec. 5.2) focuses on two rotational motions documented in experiments [65], each yielding an ensemble of $256$ scalar fields modeling the electron density of the corresponding configurations. The second case study (Sec. 5.3) investigates the $48$ intrinsic vibration modes of the above isomers, yielding $192$ ensembles of $21$ scalar fields each. In both cases, our analysis based on the topological description of the electron density corroborates documented observations [14, 15, 65]. Moreover, the fine scale of our approach (at the electronic level) enables a qualitative interpretation of the geometric conditions favoring the appareance of H-bond paths, resulting in refined suggestions for the geometrical criteria for H-bond detection in molecular graphs. We provide as additional material our C++ implementation, along with our electron density database ($4544$ datasets, about $800$ gigabytes), which may constitute a benchmark for future methods dealing with ensemble chemistry data.
0.1 Related work
This section reviews the literature related to our work, which can be classified into the following main categories.
Computational Quantum Chemistry: Quantum chemistry (QC) explores the properties of molecular systems at the granularity of the interactions between electrons and nuclei. Density functional theory (DFT) [42] is a prominent quantum mechanical framework, based on a probabilistic representation of the distribution of electrons in a molecular system, called the electron density (noted $\rho$). Based on this, Bader introduced the Quantum Theory of Atoms in Molecules (QTAIM) [3, 50], which establishes a relation between the critical points of the electron density and various chemical concepts (Sec. 1.2). QTAIM plays a central role in modern quantum chemistry as it enables a fine-scale interpretation and understanding of many experimental observations.
Molecular data analysis and visualization: In recent years, the visualization community has shown a substantial interest in molecular data (as assessed by several surveys [44]), as its complexity presents unique challenges for rendering, exploration and analysis. Several popular software platforms have emerged [40, 73] for the visualization and analysis of molecular systems using ball-and-stick representations. In particular, dedicated rendering techniques have been proposed for the interactive display of these graphs for large-scale molecular systems [12]. Moreover, for dynamic contexts, specialized techniques have also been documented for the tracking of consistent internal frames for these graphs [74, 72]. For even larger systems, molecular surfaces [16] constitute an alternate representation, based on an implicit surface derived from the molecule specification, which facilitates the visualization and analysis of large-scale phenomena (e.g., protein docking [11, 54], cavities [45, 59], pockets [46, 88] and bonding [47]). Specialized techniques for their efficient rendering have also been documented [58, 80]. While they provide a useful approximation of the shape of molecular systems, molecular surfaces only model their geometry, without accounting for the quantum mechanisms governing their structure. For a fine-scale analysis of these mechanisms, several techniques have been established on top of specific scalar descriptors (e.g., the electron density). For instance, approaches based on continuous scatterplots [69] or Voronoi segmentations [49] have been proposed for charge transfer analysis.
Topological Data Analysis: Topological methods [19, 90] have received a considerable attention by the visualization community over the last decades, given their ability to robustly extract structural patterns within scalar data [41]. They introduce a toolbox of data descriptors (persistence diagrams [20, 29], merge and contour trees [10, 27], Reeb graphs [6, 28], or Morse-Smale complexes [66, 70, 35]), where each descriptor captures specific structural relations between the critical points [51] of the input scalar field. The versatility of this toolbox has been demonstrated by the variety of domains where successful applications have been documented, including turbulent combustion [9, 32], material sciences [36, 75], fluid dynamics [43, 55], bioimaging [8], or astrophysics [71]. In particular, in quantum chemistry, these methods enable a robust implementation of QTAIM [5, 79, 57] (with tailored molecular descriptors [37]). At a conceptual level, our work deals with the study of the variability of a specific topological descriptor (the extremum graph [17], Sec. 3), within an ensemble of scalar fields, which is a topic which also received a significant attention in recent years. For instance, Athawale et al. [2] introduce a framework for the pointwise analysis of the geometrical variability of separatrices in the 2D Morse complex, based on an overlap-driven matching of mandatory critical points [31]. Also, general-purpose metrics [53] and tracking algorithms [18] have been introduced for extremum graphs. In contrast to the above methods, our work focuses on studying the global structural variability of separatrices in a specific application context (ensembles of 3D electron density fields), which comes with domain specific hypotheses, which our work specifically exploits, resulting in a simple and effective algorithm.
0.2 Contributions
This paper makes the following new contributions:
  1. 1.
    Stability measure: We present a stability measure, called bond occurrence rate, which associates to each bond path present at the equilibrium state a rate of occurrence within an input ensemble of electron densities. This measure enables the identification of densities lacking certain bond paths, enabling detailed investigations.
  2. 2.
    Algorithm: We introduce an algorithm for the computation of the above measure. It exploits the specific hypotheses of molecular systems, yielding a simple and effective computation based on a tailored, geometry-aware partial isomorphism estimation between the extremum graphs of the considered electron densities.
  3. 3.
    Case studies: We present two case studies, on rotational and vibrational dynamic effects respectively, for prominent isomers of water hexamers (the so-called Ring, Book, Cage, Prism). Our stability measure enables the identification of geometrical configurations favoring the disappearance of H-bond paths, corroborating previously documented observations. We also provide refined geometrical criteria for the emergence of H-bond paths.
  4. 4.
    Implementation: We provide a C++ implementation of our approach that can be used for reproducibility purposes.
  5. 5.
    Database: We contribute the database of electron densities generated for our case studies, i.e., 4544 datasets (about 800 gigabytes), organized in two ensembles of rotational effects (256 members each) and, for each considered isomer, 48 ensembles of vibrational effects (21 members each). This database may constitute a benchmark for future research in ensemble chemistry data.
1 Background
This section provides some technical background in chemistry and topology. Specifically, it reviews the necessary notions of molecular chemistry with an introduction to QTAIM [50]. Also, it presents a brief overview of the concepts from Topological Data Analysis which are required for the robust implementation of QTAIM. For further readings, we refer the reader to textbooks on computational topology [19, 90].
1.1 Hydrogen bonds
The term hydrogen bonds (or H-bonds) in chemistry refers to subtle yet essential attractive non-covalent interactions in molecular systems, governing their structures, properties, and reactivity.
[…]
Figure 1: Main concepts from the Quantum Theory of Atoms in Molecule (QTAIM) on a 2D slice of the electron density for a water molecule. Left: Maxima (red spheres) of the electron density (gray gradient) represent atoms, while saddles (white spheres) represent bond critical points. Arrows illustrate extreme vibrational displacement vectors (mode 00, see Sec. 2.3). Right: Bond critical points are connected to atoms via bond paths (black cylinders). In QTAIM, two atoms that are considered chemically bonded are linked by a bond path [50]. Each atom is associated with its basin (colored regions, pink, light gray and dark gray), whose boundary forms an interatomic surface (red curves on this 2D slice).
Figure 2: Multi-scale Topological Data Analysis (from left to right) with discrete Morse theory (DMT). Top: Discrete gradient fields (black: vertex-edge vectors, red: edge-triangle vectors) identify the simplices involved in no discrete vectors as critical simplices (red: minima, white: saddles). The field is progressively simplified from left to right by v-path reversal (the vectors highlighted with an ellipse have been reversed). Bottom-Left: The Morse complex decomposes the domain into stable sets. The white, dark gray and red regions only contain v-paths involved in the stable sets of minima. The light gray region only contains v-paths involved in the stable set of saddles. The unstable sets of the saddles are reported with cylinders (red and black). As the gradient field is progressively simplified (from left to right, for instance by topological persistence [20]), the resulting Morse complex is also progressively simplified.
1.2 Quantum Theory of Atoms in Molecules (QTAIM)
Quantum Theory of Atoms in Molecules (QTAIM) was initially introduced by Bader in the sixties and continuously developed since [3, 50], as it became a prominent methodology for analyzing and explaining specific phenomena which proved challenging to explain with alternative models. Drawing from the theory of gradient dynamical systems, QTAIM relies on the analysis of the electron density, noted $\rho$, which, given a molecular system, provides a three-dimensional distribution of the electronic charge in the system (Sec. 0.1). Given the specification of a molecular system, this quantity can be computed based on density functional theory (DFT) [42]. In particular, QTAIM provides a topological specification of the entities involved in molecular interactions, illustrated in Fig. 1. Local maxima of $\rho$ coincide with the nuclei of the atoms (red spheres, Fig. 1). The basin of a given atom A is defined as the region of 3D space for which all points induce a gradient integral line (i.e., a curve everywhere tangential to the gradient of $\rho$, noted $\nabla\rho$) converging to A. The set of basins forms a partition of the 3D space. Specifically, the two-dimensional boundaries between basins form zero-flux surfaces (with regard to $\nabla\rho$), called interatomic surfaces. They describe interfaces in space across which there is no net flow of electrons (red curves, Fig. 1). Local maxima of $\rho$ restricted to the interatomic surfaces correspond to saddle points [51] of $\rho$ in 3D, called bond critical points [50] (white spheres, Fig. 1). Such points are the local minima of bond paths (black cylinders, Fig. 1) connecting two nuclei, which are formed by a pair of gradient integral lines started at a saddle and converging to the corresponding maxima. According to QTAIM, if two atoms are chemically bonded (through a covalent or non-covalent bond, such as an H-bond), they have their nuclei linked by a bond path [50] and the set of bond paths form the molecular graph.
1.3 Topological Data Analysis
This section describes the necessary concepts from computational topology [19, 90] for a robust implementation of QTAIM in practice.
Input data: The input data is typically provided by the DFT computation as a scalar field $\rho:\mathcal{C}\rightarrow\mathbb{R}$, valued on the vertices of a cubical complex $\mathcal{C}$ (i.e., a three-dimensional regular grid). In practice, for implementation genericity purposes, $\mathcal{C}$ is often triangulated (via an on-the-fly emulation [79]) into a simplicial complex $\mathcal{K}$ (typically, with the Freudenthal triangulation [38]). This yields a piecewise linear (PL) scalar field $\rho:\mathcal{K}\rightarrow\mathbb{R}$, valued on the vertices of $\mathcal{K}$ and linearly interpolated on the simplices of higher dimensions. At this stage, the input $\rho$ is assumed to be injective on the vertices of $\mathcal{K}$, which is enforced in practice via symbolic perturbation [21].
Discrete gradient field: Given the above input formalization, one can leverage discrete Morse theory (DMT) [23] to design robust, combinatorial algorithms for Topological Data Analysis (as illustrated in Fig. 2). For this, the notion of a discrete gradient field can be introduced as follows. Let $\{\sigma_{i}<\sigma_{i+1}\}$ be a discrete vector. It is formed by a pair of simplices of $\mathcal{K}$, such that $\sigma_{i}$ is an $i$-simplex and $\sigma_{i+1}$ is a cofacet of $\sigma_{i}$ (see the vertex-edge vectors in black and the edge-triangle vectors in red in Fig. 2, top). $\sigma_{i+1}$ is usually referred to as the head of the vector, while $\sigma_{i}$ is its tail. A discrete vector field on $\mathcal{K}$ is defined as a collection of pairs $\{\sigma_{i}<\sigma_{i+1}\}$, such that each simplex of $\mathcal{K}$ is included in at most one discrete vector. A simplex involved in no discrete vector is called a critical simplex (larger spheres in Fig. 2). The notion of a v-path introduces a discrete analog to the concept of integral line. It is defined as a sequence of $k$ discrete vectors $\big{\{}\{\sigma_{i}^{1}<\sigma_{i+1}^{1}\},\dots,\big{\{}\{\sigma_{i}^{k}<\sigma_{i+1}^{k}\}\big{\}}$ such that (i) $\sigma_{i}^{j}\neq\sigma_{i}^{j+1}$ (i.e., the tails of two consecutive vectors are distinct) and (ii) $\sigma_{i}^{j+1}<\sigma_{i+1}^{j}$ (i.e., the tail of a vector in the sequence is a face of the head of the previous vector in the sequence) for any $1<j<k$. A v-path is said to terminate at a critical simplex $\sigma_{i}$ if $\sigma_{i}$ is a facet of the head of its last vector. By symmetry, a v-path is said to start at a critical simplex $\sigma_{i+1}$ if $\sigma_{i+1}$ is a cofacet of the tail of its first vector. We call the stable set of a critical simplex $\sigma_{i}$ the collection of v-paths which terminate in $\sigma_{i}$ (Fig. 2). By symmetry, we call the unstable set of $\sigma_{i+1}$ the collection of v-paths which start in $\sigma_{i+1}$ (cylinders in Fig. 2, bottom). Finally, a discrete vector field for which all v-paths are loop-free is called a discrete gradient field.
Several algorithms have been documented for the computation of such a discrete gradient field from an input scalar field [66, 70]. In that context, the critical simplices are discrete analogs to the notion of critical points from the smooth setting [51], with their dimension $i$ matching with the smooth notion of critical indices.
Morse complex: Given a discrete gradient field, the Morse complex is the complex formed by the collection of stable sets (Fig. 2, bottom). In practice, to cope with noise or numerical inaccuracies, the Morse complex is often simplified, either in a post-process [33] or in a pre-process [30], by modifying the discrete gradient field prior to computing the Morse complex. In particular, a pair of critical simplices $\sigma_{i}$ and $\sigma_{i+1}$ can be simplified together if there exists a unique v-path connecting them, by simply reversing the discrete vectors along the v-path [23]. Globally, the Morse complex can be progressively simplified, by iteratively reverting v-paths connecting pairs of critical simplices (Fig. 2, from left to right). In practice, several criteria can be considered for selecting a v-path for reversal. For instance, topological persistence [20, 19] is an established, general-purpose importance measure, that is well studied in Topological Data Analysis. Overall, this combinatorial approach enables a robust and straightforward implementation of QTAIM [5, 79, 57, 7]. Specifically, since in DMT discrete vectors point down [23], the QTAIM basins correspond to the stable sets of the minima of the opposite of the electron density, and the interatomic surfaces correspond to the stable sets of its $1$-saddles. Also, the bond paths are given by the unstable sets of $1$-saddles of this opposite density.
2 Electron Density Ensemble Generation
This section describes the generation of the ensemble datasets considered in our case studies (Sec. 5.2 and Sec. 5.3). These ensembles model, for selected water clusters (Sec. 2.1), proton tunneling (Sec. 2.2) and molecular vibrations (Sec. 2.3). These are two important phenomena naturally occurring in chemical processes, which can trigger isomerization – transforming a molecule into another, with identical composition, but with a distinct spatial arrangement of atoms, resulting in distinct chemical properties. Isomerization plays a crucial role in many areas of chemistry (e.g., drug design) and understanding the conditions favoring its occurrence is an important challenge in quantum chemistry.
Figure 3: Extremum graphs of opposite electron density ($\rho^{\prime}\in(-126,0)$), at various stages of bond graph extraction, for a Prism water hexamer (transparent: isosurface for $\rho^{\prime}=-0.1$). (a) The presence of spurious minima (red sphere) requires topological simplification [48]. Minima are colored by topological persistence (Sec. 1.3). In this example, the spurious minimum (red) has a persistence of $4\times 10^{-4}$, while the first non-spurious one has a persistence of $3.6\times 10^{-2}$ (two order of magnitude gap). (b) The presence of spurious pairs of $1$ and $2$-saddles (orange and black spheres respectively) requires a step of simplification, achieved by v-path reversals, ordered by function value difference. Here, the last spurious saddle-saddle v-path exhibits a function difference of $4\times 10^{-5}$, while the first non-spurious one has a difference of $5\times 10^{-3}$ (two order of magnitude gap). (c) The simplified extremum graph is represented with its nodes and arcs colored by $\rho^{\prime}$ (nodes: value at the minimum, arcs: value at the saddle), enabling a clear distinction between oxygen (red) and hydrogen (pink) atoms as well as between covalent (pink) and non-covalent (white to black) bonds. (d) The resulting bond graph represents the identified atoms (red: oxygen, white: hydrogen) and bond types (dark gray: covalent, light gray: H-bond).
2.1 Water cluster selection
The study of water clusters, (H2O)${}_{n_{m}}$, is essential for the understanding of the structure of liquid water and ice [85, 25]. It provides key insights into the mechanisms of many processes, such as water aggregation, microsolvation [67, 39, 87] or proton transfer [83].
For a given number of molecules $n_{m}$, water clusters can exhibit distinct isomeric forms (i.e., geometrical arrangements), as each water molecule can partcipate in four H-bonds. Each isomer has a specific hydrogen bond network (i.e., set of H-bonds), resulting in unique chemical properties. Moreover, isomers with comparable energies can easily rearrange into one another through a process called interconversion [24]. In our work, we focus on these relatively simple yet chemically important systems due to their extensive documentation and ubiquity in chemistry. For example, several studies document proton tunneling (Sec. 2.2) or molecular vibration (Sec. 2.3) for these clusters, hence helping validating our approach. Also, their geometric scale, between micro- and macro-systems, typically makes them good candidates in chemistry for investigating novel models or experimental protocoles, before considering more complex systems (see Sec. 6).
Specifically, we focus on water hexamers ($n_{m}=6$), which are prominent examples of water clusters [25], as they constitute the smallest three-dimensional cluster whose non-quasiplanar isomers (the Book, Cage and Prism) exhibit lower energies than quasiplanar isomers (such as the Ring). For this reason, (H2O)6 is often referred to in the literature as “the smallest drop of water” [56]. High resolution spectroscopy as well as state-of-the-art quantum calculations [68, 61] indicate that the Prism, the Cage and the Book are the isomers of (H2O)6 with the lowest energies. Consequently, they likely coexist, even at low temperatures. Therefore, we focus our study on these three isomers as well as the Ring isomer, which is the quasiplanar structure with the lowest energy.
2.2 Tunneling pathways
To evaluate the chemical relevance of our approach, we first study the Prism isomer under the effect of proton tunneling - a phenomenon in which a molecule in its equilibrium state rearranges to an equivalent (i.e., isoenergetic) structure, for example by permuting equivalent atoms [62]. Proton tunneling in the Prism isomer of (H2O)6 was observed experimentally and confirmed by calculations [65, 81, 14]. Specifically, far-infrared vibration-rotation tunneling (FIR-VRT) spectroscopy observations report characteristic splitting patterns. These can be explained by two distinct tunneling mechanisms, associated to specific rotations of water molecules: (i) An anti-geared rotation, breaking one H-bond, and (ii) a geared motion, simultaneously breaking two H-bonds. Each of these two motions can be modeled by a specific pathway (i.e., a temporal sequence of transitory structures between the two equivalent states), which can be estimated with the ring-polymer molecular dynamics (RPMD) approach [64]. For each step $\alpha_{k}\in(0,255)$ of the pathways reported in [65], the electron density was computed with a non-relativistic DFT method [42], with the PBE0 functional [86] and the TZ2P basis, as implemented in the ADF software [77] and facilitated through PyADF[22]. The computations were carried out on a $256^{3}$ regular grid, with a sample spacing of $0.05$ Angstrom (Å).
2.3 Molecular vibrations
Molecular vibrations are intrinsic behaviors of molecular systems. They describe the relative motion of atoms (occurring at all temperatures) which does not change the position of the molecular center of mass. In the following, although our molecular vibration modeling follows the typical approach in chemistry [82], we still specify its details in order to precisely document our data generation process, thereby enabling reproducibility. A system composed of $n_{a}$ atoms is a subject to $n_{m}=3n_{a}-k$ vibrational degrees of freedom, with $k=3+3$ in non-linear molecules (to discard the 3 translational and 3 rotational degrees of freedom of the molecular center of mass). In particular, for a water hexamer, (H2O)6, $n_{a}=18$ and $n_{m}=48$. The complex motion induced by these degrees of freedom can be decomposed into simpler components, referred to as the normal modes of vibrations. Each mode describes the motion of all atoms occurring with a specific frequency and phase.
The equilibrium structure of each isomer, noted ${\mathcal{R}}^{0}$, was optimized with the PBE0 TZ2P model (Sec. 2.2) and the harmonic vibrational modes, noted ${\mathcal{Q}}$, were computed via normal mode analysis (NMA), a standard method for studying molecular vibrations under the harmonic oscillator approximation [82]. The displacement of a nucleus $i$ along each of the $n_{m}$ vibrational normal vectors ${\mathcal{Q}}_{j}\in\mathbb{R}^{3n_{a}}$ is then given by:
${\Delta}_{i,j}=s_{i}{\mathcal{L}}_{i,j}{\mathcal{Q}}_{j},\qquad i\in\{1,\ldots,n_{a}\},$
where $s_{i}$ is a coefficient accounting for the masses of the oscillating nuclei and where $\mathcal{L}_{i,j}$ is constructed from the second-order (Cartesian) derivatives of the potential energy at the equilibrium atomic positions.
Next, let $\alpha_{k}$ be an integer in the range $(0,20)$, which constitutes a sufficient sampling for our case study (to enable the observation of clear transitions in H-bond configurations, see Sec. 5.3). The position of a nucleus $i$ in the perturbed geometry $k$ of the vibration mode $j$ is:
$\mathcal{R}_{i,j,k}=\mathcal{R}^{0}_{i}+0.4\cdot(\alpha_{k}-10)\cdot\mathcal{A}_{h}\cdot\Delta_{i,j},$
where $\mathcal{A}_{h}$ is a classical estimation of the turning point of the harmonic oscillatory motion [26]. Fig. 1 illustrates the resulting vectors on a water molecule at mode $j=0$, for extreme displacements (i.e., $k\in\{0,20\}$).
Finally, given the obtained out-of-equilibrium structures $\mathcal{R}_{j,k}$, we considered $n_{m}=48$ ensembles of electron densities for each selected isomer (Sec. 2.1), calculated with the the same computational model as used in Sec. 2.2. Each ensemble describes one vibrational normal mode and counts $n_{v}=21$ electron densities generated from: (i) the isomer at equilibrium ($\alpha_{10}$) and (ii) its $20$ derived geometries, obtained by displacing each atom along its vibrational normal vector ($10$ in negative directions $\alpha_{k}\in\{0,9\}$, and $10$ in positive directions $\alpha_{k}\in\{11,20\}$).
3 Bond Graph Extraction
This section documents the topological analysis of a single electron density dataset, for the automatic extraction of a graph, which we call bond graph, representing the covalent and non-covalent bonds in the system. While the pipeline described in this section is typical of related work [5, 57], we emphasize here the details that are specific to both water hexamers and the analysis of the stability of their hydrogen bonds.
3.1 Chemical features of interest specification
Given the electron density $\rho:\mathcal{K}\rightarrow\mathbb{R}$ of a water hexamer, and following QTAIM (detailed in Sec. 1.2), the goal of our analysis is to extract a geometrical representation of the following features of interest: oxygen atoms, hydrogen atoms, covalent bonds, H-bonds.
As described in Sec. 1.3, since in discrete Morse theory discrete vectors are pointing down, we will consider in the following the opposite electron density, i.e., the field $\rho^{\prime}:\mathcal{K}\rightarrow\mathbb{R}$, such that $\rho^{\prime}(v)=-\rho(v)$ for all the vertices $v\in\mathcal{K}$. Then, the nucleus of each atom can be identified as a local minimum of $\rho^{\prime}(v)$. Within each molecule, a covalent bond can be extracted as the bond path (i.e., the unstable set, Sec. 1.3) of a bond critical point (i.e., a $1$-saddle of $\rho^{\prime}(v)$ located within a molecule). Finally, H-bonds can be extracted as the remaining bond paths which connect a hydrogen atom to an oxygen atom. We describe in the following two sections how to robustly extract these features.
3.2 Topological characterization
This section documents the topological analysis of the input opposite electron density field $\rho^{\prime}:\mathcal{K}\rightarrow\mathbb{R}$.
Scalar data is often impacted by noise, even with smooth data such as the electron density, as shown in Fig. 3(a), where spurious minima occur in $\rho^{\prime}$, due to local numerical inaccuracies. To address this issue, we pre-process $\rho^{\prime}$ with persistence-driven simplification. Specifically, the minima associated with a measure of topological persistence [20, 19] lower than a threshold $\epsilon$ are simplified with local perturbations of the scalar data, using the algorithm presented by Lukasczyk et al. [48]. In particular, $\epsilon$ is selected such that only $18$ minima remain in the field ($3$ atoms per water molecule, for each of the $6$ molecules of the system). In practice, $1\times 10^{-3}$ is a typical cut-value for $\epsilon$ (see Fig. 3).
Figure 4: Instability of unstable sets on a 2D Gaussian mixture (with negative coefficients, gray gradient; four isolines are shown, from black to red). From (a) to (d), the minimum $A$ (red sphere) is translated along a direction parallel to the axis $B-C$ (red arrow). From (b) to (c), only a slight displacement of $A$ is sufficient to switch the position of the saddle $D$ (white sphere), resulting in a drastic change in the structure of the unstable sets: $A$ is now connected to $C$ instead of $B$. Note that this instability is not necessarily correlated to topological persistence (radius of the spheres, see Sec. 1.3), as the saddle $D$ is involved in a persistence pair more persistent than that of $E$. This example further motivates our investigation of the practical stability of QTAIM bond-paths.
Similarly, numerical inaccuracies can also generate spurious pairs of saddles in the input field $\rho^{\prime}$ (Fig. 3(b)). As described in Sec. 1.3, these can be removed via iterated v-path reversals. Specifically, we revert a v-path connecting a $2$-saddle down to a $1$-saddle if their function value difference is less than $1\times 10^{-4}$.
Once the above two pre-processes are completed (to remove spurious minima and $1$-saddles), the extremum graph [17] is computed by collecting the unstable sets (Sec. 1.3) of each remaining $1$-saddle. Formally, the extremum graph of $\rho^{\prime}$, noted $\mathcal{E}(\rho^{\prime})$, is a graph with a set of nodes $\mathcal{N}(\rho^{\prime})$ and a set of arcs $\mathcal{A}(\rho^{\prime})$, such that each node $n\in\mathcal{N}(\rho^{\prime})$ represents exactly one minimum of $\rho^{\prime}$, and each arc $a\in\mathcal{A}(\rho^{\prime})$ represents the unstable set of a $1$-saddle of $\rho^{\prime}$ linking two of its minima.
Note that, for a reliable structural analysis of the extremum graph, the geometrical accuracy of the unstable sets is of paramount importance. For that, among the different discrete gradient algorithms studied in the literature [66, 70, 63, 34, 35, 78], we use the approach by Gyulassy et al. [34], as it provides improved accuracy and provable convergence to the underlying continuous flow (for its stochastic variant).
Interestingly, from a mathematical point of view, as their name suggests, unstable sets are fundamentally unstable constructions, even when considering geometrically accurate discrete gradient fields. Fig. 4 shows an example of a 2D scalar field, where an arbitrarily small perturbation results in a drastically different extremum graph. Despite this instability, unstable sets are widely used in QTAIM for characterizing covalent and non-covalent bond paths (Sec. 1.2). This paradox further motivates our ensemble-based H-bond stability analysis.
3.3 From topological to chemical features
This section documents the extraction of the bond graph from the extremum graph computed in the previous section.
Since oxygen and hydrogen atoms have notably different atomic masses, these can be easily distinguished based on their opposite electron density values ($\rho^{\prime}$), typically in our database around $-8$ for oxygen atoms and above $-1$ for hydrogen atoms. Thus, we classify as oxygen atoms all the atoms with an opposite electron density value below $-4$, and as hydrogen atoms all the atoms above this value.
Similarly, since they involve strong electrostatic interactions, covalent bonds can be easily distinguished from non-covalent bonds based on the opposite electron density value ($\rho^{\prime}$) of their corresponding bond critical point. Specifically, we classify as covalent bond an unstable set from the extremum graph which connects an oxygen atom to a hydrogen atom, and whose bond critical point has an opposite electron density value below $-0.1$. Symmetrically, unstable sets connecting an oxygen atom to a hydrogen atom, and whose bond critical point has an opposite electron density value above $-0.1$ are classified as non-covalent bonds. Note that this isovalue of opposite electron density ($-0.1$) is often selected for representing the geometry of a molecule via isosurfacing (Fig. 3). In certain cases, unstable sets may connect two oxygen atoms, which does not characterize a priori any chemically relevant interaction in this context. We will refer to such connections as misconnected H-bonds (see Sec. 5.1).
Finally, the bond graph is obtained by collecting the covalent and hydrogen bonds identified above, see Fig. 3(d). This graph will be later exploited in our use cases (Sec. 5) to evaluate bond stability.
Figure 5: Geometry-aware partial isomorphism estimation between the extremum graphs of the Prism water hexamer at equilibrium (top) and under large vibrational displacement (bottom). (a) First, an optimal assignment matches the nodes together (colors), in order to minimize the sum of their 3D distances (Eq. 1). (b) This matching is extended to the arcs (colors) by matching arcs together if their nodes are matched by the optimal node assignment, possibly yielding a partial isomorphism (unmatched arcs are shown in white). (c) For this ensemble (only two graphs), the bond occurrence rate reports a value of $1$ (red) for all arcs involved in the partial isomorphism and of $0.5$ (black) for the others.
Figure 6: Bond occurrence analysis for the Prism (vibration mode 00). (a) Extremum graphs for the input vibrational displacements (black to red: negative to positive translations, equilibrium: white). (b) Partial isomorphisms to the equilibrium state (colored arcs, white: unmatched arcs). (c) Bond occurrence rate (black to red). (d-h) Bond occurrence rate for various configurations, as a function of the vibrational displacement: (d) missing H-bond ($\alpha_{0}$), (e) misconnected H-bonds ($[\alpha_{1},\alpha_{2}]$), (f) valid H-bonds ($[\alpha_{3},\alpha_{12}]$), (g) misconnected H-bonds ($[\alpha_{13},\alpha_{17}]$), (h) missing H-bonds ($[\alpha_{18},\alpha_{20}]$). Valid H-bonds occur around the equilibrium, while misconnected and missing H-bonds (arrows) occur for intermediate and large displacements.
4 Analysis of Ensembles of Extremum Graphs
This section describes our approach for estimating a geometry-aware partial isomorphism between two extremum graphs, representing two distinct electron densities. This partial isomorphism is later exploited in the definition of our stability measure (called bond occurrence rate).
4.1 Overview
Let $\rho^{\prime}_{i}:\mathcal{K}\rightarrow\mathbb{R}$ and $\rho^{\prime}_{j}:\mathcal{K}\rightarrow\mathbb{R}$ be two opposite electron densities representing a given molecular system under distinct geometrical configurations (e.g., distinct vibrational displacements). Moreover, let $\mathcal{E}(\rho^{\prime}_{i})$ and $\mathcal{E}(\rho^{\prime}_{j})$ be their respective extremum graphs, with their node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$ and their arc sets $\mathcal{A}(\rho^{\prime}_{i})$ and $\mathcal{A}(\rho^{\prime}_{j})$, see Sec. 3.2.
The goal of the approach documented in this section is to establish a reliable correspondence between these two extremum graphs. For this, we first establish a correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ between the nodes of the extremum graphs (see Sec. 4.2 and Fig. 5(a)), prior to extending it to their arcs, in the form of an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}:\mathcal{A}(\rho^{\prime}_{i})\rightarrow\mathcal{A}(\rho^{\prime}_{j})$ (see Sec. 4.3 and Fig. 5(b)). To do so, we specifically exploit the fact that our database models water hexamers. Therefore, at this stage, each extremum graph counts exactly $18$ nodes (representing the $3$ atoms of the $6$ water molecules) and we have $|\mathcal{N}(\rho^{\prime}_{i})|=|\mathcal{N}(\rho^{\prime}_{j})|$, which is a valuable simplifying hypothesis.
4.2 Optimal node assignment
While the original molecule specifications (i.e., the atom locations) are known for both opposite densities $\rho^{\prime}_{i}$ and $\rho^{\prime}_{j}$ (this is the input of the DFT computation, Sec. 2), in practice, the locations of their resulting minima do not exactly coincide to this specification due to both spatial quantization artifacts and the imbalance between the atomic masses of the oxygen and hydrogen atoms (which displaces the minima, especially those associated to hydrogen atoms). Thus, a robust algorithm needs to be considered to establish a reliable correspodence between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. For this, we consider the following balanced assignment problem (the simplest form of optimal transport [60]).
Let $\phi_{i\rightarrow j}:\mathcal{N}(\rho^{\prime}_{i})\rightarrow\mathcal{N}(\rho^{\prime}_{j})$ be a bijection between the node sets $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. Next, let $E(\phi_{i\rightarrow j})$ be the following energy term:
$E(\phi_{i\rightarrow j})=\sum_{n\in\mathcal{N}(\rho^{\prime}_{i})}||n-\phi_{i\rightarrow j}(n)||_{2}.$ (1)
The term $E(\phi_{i\rightarrow j})$ simply evaluates the sum of the distances in 3D space between the minima of $\rho^{\prime}_{i}$ and their images by $\phi_{i\rightarrow j}$ (intuitively, the sum of the geometrical mismatches between minima). Finally, we select as correspondence $\phi_{\mathcal{N}_{i\rightarrow j}}$ a global minimizer of $E(\phi_{i\rightarrow j})$:
$\phi_{\mathcal{N}_{i\rightarrow j}}=\operatorname*{arg\,min}_{\phi_{i\rightarrow j}\in\Phi_{i\rightarrow j}}E(\phi_{i\rightarrow j}),$
where $\Phi_{i\rightarrow j}$ denotes the set of all bijections between $\mathcal{N}(\rho^{\prime}_{i})$ and $\mathcal{N}(\rho^{\prime}_{j})$. In practice, the solution of this assignment problem, Fig. 5(a), can be easily obtained with either exact [52] or approximate [4] algorithms.
4.3 Induced partial isomorphism
The optimal node assignment computed in the previous section is extended into an induced partial isomorphism $\phi_{\mathcal{A}_{i\rightarrow j}}$ as follows. Let $a_{i}\in\mathcal{A}(\rho^{\prime}_{i})$ and $a_{j}\in\mathcal{A}(\rho^{\prime}_{j})$ be two arcs of $\mathcal{E}(\rho^{\prime}_{i})$ and $\mathcal{E}(\rho^{\prime}_{j})$ respectively. Next, we assign $a_{i}$ to $a_{j}$, i.e., $\phi_{\mathcal{A}_{i\rightarrow j}}(a_{i})=a_{j}$, if and only if:
$\phi_{\mathcal{A}_{i\rightarrow j}}(a_{i})=a_{j}\Longleftrightarrow\begin{cases}\phi_{\mathcal{N}_{i\rightarrow j}}(a_{i}^{1})=a_{j}^{k}\\ \phi_{\mathcal{N}_{i\rightarrow j}}(a_{i}^{2})=a_{j}^{l}\end{cases},$
where $a_{i}^{1}$ stands for the first node of $a_{i}$ and $a_{i}^{2}$ for the second, and where $(k,l)\in\{(1,2),(2,1)\}$. In other words, an arc $a_{i}$ will be assigned to an arc $a_{j}$ if the nodes of $a_{i}$ are mapped through $\phi_{\mathcal{N}_{i\rightarrow j}}$ to those of $a_{j}$.
If the resulting map $\phi_{\mathcal{A}_{i\rightarrow j}}:\mathcal{A}(\rho^{\prime}_{i})\rightarrow\mathcal{A}(\rho^{\prime}_{j})$ is a bijection, we will say that it describes a complete isomorphism between $\mathcal{E}(\rho^{\prime}_{i})$ and $\mathcal{E}(\rho^{\prime}_{j})$. Otherwise, it only describes a partial isomorphism, e.g., Fig. 5(b).
Figure 7: H-bond path configurations for the Prism (vibration mode 00), according to their unstable sets (white: saddles, red: minima, black curves: unstable sets). A 2D slice of the electron density is shown in the background with three isolines. The corresponding partial isomorphisms are shown with colors in the insets. (a) At equilibrium ($\alpha_{10}$), the unstable set of the saddle $A$ connects it to hydrogen (left) and oxygen (right) atoms. (b) Under intermediate displacement ($\alpha_{2}$), the saddle $A$ increases its $\rho^{\prime}$ value (white isoline). The geometry of its unstable set is altered such that it now connects it to two oxygen atoms. We refer to this configuration as a misconnected H-bond. (c) Under large displacement ($\alpha_{0}$), the saddle $A$ disappears. The electron density field has been sufficiently modified by the vibrational displacement to cancel the corresponding saddle-saddle persistence pair [13]. We refer to this configuration as a missing H-bond.
4.4 Bond occurrence rate
In this section, we leverage the partial isomorphism introduced above to provide an estimation of the stability of an H-bond path.
Let $\{\mathcal{E}(\rho^{\prime}_{1}),\mathcal{E}(\rho^{\prime}_{2}),\dots,\mathcal{E}(\rho^{\prime}_{n})\}$ be the set of $n$ extremum graphs generated with the approach described in Sec. 3.2, from an ensemble of $n$ electron densities (generated as documented in Sec. 2).
Given an arc $a$ from an extremum graph $\mathcal{E}(\rho^{\prime}_{i})$ (with $i\in[1,n]$), we estimate its bond occurrence rate $\mathcal{O}(a)$ as follows:
$\mathcal{O}(a)={{1}\over{n}}\sum_{j\in[1,n]}|\{\phi_{\mathcal{A}_{i\rightarrow j}}(a)\leavevmode\nobreak\ |\leavevmode\nobreak\ \leavevmode\nobreak\ \phi_{\mathcal{A}_{i\rightarrow j}}(a)\neq\emptyset\}|.$
In other words, the bond occurrence rate $\mathcal{O}(a)$ measures the number of times a bond $a$ from an extremum graph $\mathcal{E}(\rho^{\prime}_{i})$ has been matched to another bond in the ensemble (divided by the size of the ensemble). This quantity, Fig. 5(c), is expected to equal $1$ for stable bonds (e.g., covalent bonds) and to be in the range $[{{1}\over{n}},1]$ for unstable bonds.
5 Results
This section presents experimental results obtained with our approach, implemented in C++ within TTK [79, 7]. Our experimental data Sec. 2 is available at this address: https://github.com/thom-dani/BondMatcher. In both case studies (Secs. 5.2 and 5.3), our bond occurrence rate (Sec. 4.4) played an instrumental visual role as it enabled, with only one visualization (at equilibrium), the quick identification of the H-bonds involved in transformations (instead of exploring one by one the members of the considered ensemble). This quick identification eased their selection for refined inspections, detailed in Secs. 5.2 and 5.3.
5.1 Geometrical interpretation
Running the topological data analysis pipeline described in Sec. 3 yields a database of extremum graphs. Fig. 7 presents a classification of the H-bond configurations that we have observed in this database, according to their associated unstable sets (Sec. 1.3). This figure shows that, under vibrational displacement, the bond critical point (Sec. 1.2) associated to an H-bond path gradually increases its $\rho^{\prime}$ value. Depending on the amplitude of the vibrational displacement, this can result in two degenerate configurations. First, a misconnected H-bond, Fig. 7(b), occurs when the $\rho^{\prime}$ value of the bond critical point increased to an intermediate level, such that the geometry of its unstable set is altered to the point that it connects two oxygen atoms together (a configuration which is not chemically relevant in this context). Second, a missing H-bond, Fig. 7(c), occurs when the $\rho^{\prime}$ value of the bond critical point increased so much that the corresponding saddle-saddle pair has been canceled [13]. This figure suggests that, depending on the geometric transformation induced by the vibration, the transition from a valid H-bond path at equilibrium to a missing one is gradual (via a misconnection step), which may relate to experimental studies discussing the role of oxygen-oxygen contractions on vibrations of liquid water [84].
Fig. 6 confirms this observation by inspecting the entire set of $21$ vibrational displacements for the fist vibrational mode of the Prism. Specifically, it shows that missing H-bonds occur at the extremities of the displacement spectrum ($\alpha_{0}$ and $[\alpha_{18},\alpha_{20}]$) and that misconnected H-bonds occur within immediately adjacent intervals ($[\alpha_{1},\alpha_{3}]$ and $[\alpha_{13},\alpha_{17}]$). This confirms that the transition from valid (around the equilibrium), to misconnected, to missing H-bond is indeed gradual along this vibration. Fig. 6(c) also shows the bond occurrence rate (Sec. 4.4) for these configurations. It reports lower rates for misconnected H-bonds, indicating the transitional aspect of this configuration.
In the following, we will consider that a bond identified at equilibrium with the procedure described in Sec. 3.3 is unstable if its bond occurrence rate (Sec. 4.4) is smaller than $1$. Otherwise, it is stable.
Fig. 8 investigates the relation between the unstable bonds of the Prism (relative to vibrations) and topological persistence [20], a generic importance measure often used to estimate prominence and stability in Topological Data Analysis [19]. In particular, this figure shows that, in this example, the bonds identified as unstable are defined relatively to saddles which are also involved in the least persistent saddle-saddle pairs (similar observations, not reported here, were made for the Cage isomer). This indicates that such low-persistence saddles are more likely to undergo, under vibrations, the transitions observed in Fig. 7. Still, note that in theory, instability can still be observed with unstable sets defined relative to the most persistent saddles, as illustrated in Fig. 4. Then, although persistence and bond stability appear to be correlated in our experiments, it is not necessarily the case in generality.
Finally, we conclude this section by observing that in all our experiments (either on vibrational effects or proton tunneling), the bonds known as covalent have always been reported with a bond occurrence rate of $1$. This indicates that our stability measure (Sec. 4.4), for these bonds already, is concordant with established chemical expectations.
Figure 8: Observing the persistence [20] and bond occurrence rate on the Prism (equilibrium). (a) Saddle persistence pairs (persistence: black to red) : $1$-saddles create cycles in the extremum graph, while $2$-saddles represent the topological discs which fill these cycles. (b-c) Bond occurrence rate (black to red), for mode 00 (b) and mode $2$ (c). In this example, the saddles defining unstable H-bonds (arrows) are also involved in the least persistent saddle pairs (black, gray, pink).
Figure 9: H-Bond analysis along the geared tunneling pathway on the Prism. (a) Extremum graphs for the input pathway (time: black to red). (b) Partial isomorphisms to the pathway mid-point (colored arcs, white: unmatched arcs). (c) From $t=0$ to $t=120$, the system includes, among others, two H-bonds (black and gray) respectively involving a blue and orange hydrogen atoms. (d) From $t=121$ to $t=124$, the black H-bond has disappeared and is replaced by a misconnected H-bond (thin black cylinder). (e) At $t=125$, the dark gray H-bond has disappeared and is also replaced by a misconnected H-bond. (f) From $t=126$ to $t=133$, a new H-bond has appeared (pink), involving the former, orange hydrogen atom. (g) As of $t=134$ (till $255$), a new H-bond has appeared (red), involving a green hydrogen atom. (h-j) Bond critical point density, (Ac--H) bond length and (Ac--Dn, Ac--H) bond angle over time for the black, gray, pink and red H-bonds. Our analysis captures the structural change involved in this pathway, with a clear breaking point for the two H-bonds involved in this geared tunneling motion, characterized by a sharp decrease in bond critical point densities (below $1.5\times 10^{-2}$), as well as a sudden increase in bond length (above $2.3$) and angle (above $14$°).
Figure 10: H-Bond analysis along the anti-geared tunneling pathway on the Prism. (a) Extremum graphs for the input pathway (time: black to red). (b) Partial isomorphisms to the pathway mid-point (colored arcs, white: unmatched arcs). (c) From $t=0$ to $t=115$, the system includes an H-bond (gray) involving a red hydrogen atom. (d) From $t=116$ to $t=123$, the gray H-bond has disappeared and is replaced by a misconnected H-bond (thin black cylinder). (e) As of $t=124$ (till $255$), a new H-bond has appeared (pink), involving a blue hydrogen atom. (f-h) Bond critical point density, (Ac--H) bond length and (Ac--Dn, Ac--H) bond angle over time for the gray and pink H-bonds. Our analysis captures the structural change involved in this pathway, with a clear breaking point, similar to the geared motion: bond critical point density below $1.5\times 10^{-2}$, sudden increase in bond length (above $2.3$) and angle (above $14$°).
5.2 Case Study 1: H-Bond stability in tunneling reaction
To evaluate the chemical relevance of our work, we first study the Prism under proton tunneling (Sec. 2.2), a phenomenon for which theoretical models have been validated by experimental observations [65].
Fig. 9 presents the results obtained by our approach, for the pathway describing a geared tunneling motion, where two water molecules of the Prism rotate simultaneously, resulting in a characteristic switch in the configuration of H-bonds [65]. This figure reports the geometry-aware partial isomorphisms to the pathway mid-point, as computed by our approach (Sec. 4). Specifically, the bond matching provided by our method enables an individual analysis of the geometrical properties of each bond through time. In particular, this analysis identifies two H-bonds with an occurrence rate lower than $1$ (black and gray), which successively disappear (at $t=121$, then $t=124$). Symmetrically, two new H-bonds (pink and red) successively appear (at $t=126$ and $t=134$). As shown on the companion plots, the disappearance of the black and gray H-bonds is associated with a drastic decrease in the respective bond critical point density ($\rho$), along with an abrupt increase in (Ac--H) bond length and (Ac--Dn, Ac--H) bond angle. The resulting spikes in the plots materialize a breaking point in these geometrical indicators, beyond which H-bond paths disappear. A symmetrical observation can be made for the appearing H-bonds (pink and red). These plots, enabled by our approach, confirm the brevity of the H-bond configuration switch ($14$ time steps out of $255$). They also provide limit values, for the bond critical point density ($1.5\times 10^{-2}$), bond length ($2.3$) and angle ($14$°), beyond which an H-bond path can be considered as disappearing. Fig. 10 provides a similar analysis, with identical conclusions, for an anti-geared tunneling motion, where only one H-bond disappears (and re-appears) along the pathway.
5.3 Case-Study 2: H-Bond stability in molecular vibrations
We now investigate the stability of H-bond paths under molecular vibrations, with the data generated as specified in Sec. 2.3. For this, for each considered isomer of water hexamer (the Ring, the Book, the Cage and the Prism, Sec. 2.1), we analyzed $21$ vibrational displacements under $48$ vibrational modes. Fig. 11 reports the number of H-bonds identified at equilibrium (Sec. 3.3), which have been reported as unstable (i.e., with a bond occurrence rate smaller than $1$, Sec. 4.4), as a function of the vibration mode. The first important observation is that the Ring and the Book do not report any unstable bonds. In contrast, the Cage reports few vibration modes for which unstable H-bonds are discovered, while the Prism reports more. Note that, for the latter two isomers, the analysis had to be restricted to the modes 00 to $36$, as beyond, for each mode, several electron density datasets were lacking some hydrogen atoms (i.e., they were initially exhibiting less than $18$ minima of $\rho^{\prime}$, see Sec. 5.5). Specifically, the Cage reports $2$ H-bond paths at equilibrium, which have been identified as unstable in at least one vibration mode (Fig. 12). The Prism reports $3$ unstable H-bond paths (visible in Fig. 8).
Figure 11: Number of unstable H-bond paths as a function of the vibration mode number, for the Cage (left) and the Prism (right). No unstable H-bond paths are reported by our approach for the Ring and the Book.
Fig. 13 investigates the geometrical characteristics, at equilibrium, of the different identified H-bonds (Sec. 3.3) for the $4$ isomers considered in our study. In particular, in these plots, each H-bond is represented by a point: $6$ for the Ring, $7$ for the Book, $8$ for the Cage and $9$ for the Prism. Moreover, the H-bonds reported as unstable (with an occurrence rate below $1$) are marked in gray. The purpose of this experiment is to observe if the unstability of an H-bond under molecular vibrations can be predicted based on its geometry at equilibrium, in particular in the light of the limit values identified in the previous case study (Sec. 5.2).
Fig. 13 confirms that, at equilibrium, the geometrical indicators (bond length and angle) of the H-bonds of the Ring and Book are associated to different ranges than for the Cage and Prism. This is obvious with the Ring, which has a highly symmetrical structure (BondMatcher: H-Bond Stability Analysis in Molecular Systems), leading to compact point distributions, Fig. 13(a). Also, these values are far from the limit values identified in Sec. 5.2. In contrast, the H-bonds identified as unstable in the Cage and Prism (gray points) are also the H-bonds whose geometrical indicators are the closest to the limit values identified in Sec. 5.2. This indicates that these unstable H-bonds are already, initially at equilibrium, close to the H-bond breaking points identified in the tunneling case study (Sec. 5.2), and that they are, consequently, more conducive to disappearance under molecular vibrations. As discussed in Fig. 12, such an H-bond disappearance can eventually transform an isomer into another. Then, the range of geometrical indicators for unstable H-bonds (gray spheres) – bond length: $(2,2.3)$, bond angle: $(8,14)$° – can be interpreted as uncertainty intervals, where H-bond paths are unstable and likely to disappear under vibrations.
Figure 12: Bond occurrence rate for the Cage at two vibration modes (left: mode $18$, right: mode $5$). (a,c) Extremum graphs for the input vibrational displacements (black to red). (b) At mode $18$, only one H-bond path is reported as unstable (black). (d) At mode $5$, two H-bond paths are unstable (black and gray). At mode $18$, the breaking of the unstable H-bond (b) makes the extremum graph become isomorphic to the Book.
Figure 13: Geometrical indicators, (Ac--Dn, Ac--H) bond angle as a function of (Ac--H) bond length, for the H-bonds identified at equilibrium for the (a) Ring, (b) Book, (c) Cage, and (d) Prism. The H-bonds reported as unstable in at least one vibration mode are shown in gray (red otherwise).
5.4 Computational aspects
The timings for the quantum chemistry calculations are approximately $6000$ seconds for the geometry optimization and frequency calculations for each isomer (Sec. 5.3) and $300$ seconds for each (single-point) electron density computation. Our topological analysis pipeline (Sec. 3) involves algorithms whose running time is at most quadratic, in the worst case, with the number of vertices in the grid. In practice, this pipeline required approximately $200$ seconds of computation on a desktop computer for each electron density. Once the database of electron densities is processed, a database of extremum graphs is available. Its processing by the approach described in Sec. 4 was very fast in practice, as each extremum graph is extremely small ($18$ nodes), resulting in runtimes below a second for each ensemble, even when considering the largest ensembles ($256$ members for each tunneling pathway).
5.5 Limitations
A first limitation, which prevented us from analyzing vibrations modes beyond mode $36$ for the Cage and Prism, is that our input database contains a few datasets lacking some critical points corresponding to hydrogen atoms, with a number of local minima of $\rho^{\prime}$ initially smaller than $18$. These are high-frequency modes describing the O--H stretching involving the non-H-bonded H atoms. These modes may be susceptible to effects not explicitly handled in the harmonic oscillator model. Moreover, our model artifically extends the viable range of these motion amplitudes. As a result, the displacements of these H atoms may be overestimated, such that their corresponding minima may sometimes vanish. This might be addressed by re-generating the data on a finer grid (e.g., $512^{3}$), at the expense of larger computation times. Another possibility would be to consider an unbalanced variant of the assignment problem from Sec. 4.2. Nevertheless, this would result in certain covalent bonds (involving the missing hydrogen atoms) being reported as unstable, which would challenge the overall interpretation.
In principle, large atom displacements may also challenge our optimal node assignment procedure (Sec. 4.2), in which case, further costs may need to be considered to extend the assignment energy with domain knowledge. However, we have not observed extreme displacements of minima in our experiments, despite extending the range of amplitudes of atomic motions beyond the classical harmonic model (Sec. 2.3). Our database considers intrinsic molecular vibrations that result in practice in moderate displacements of minima, such that our node matching procedure (Sec. 4.2) always provided relevant assignments.
6 Conclusion
This application paper presented BondMatcher, a framework for studying the stability of QTAIM H-bond paths in molecular systems. Specifically, by computing geometry-aware partial isomorphisms between extremum graphs extracted with discrete Morse theory (DMT), our work enables a robust tracking of H-bond paths in ensembles of electron densities. This detailed per-bond tracking enabled the identification of unstable H-bond paths, as well as their individual geometrical analysis through the ensemble. This led to the following empirical findings:
  • Despite their theoretical instability (Fig. 4), QTAIM bond paths have shown in our experiments to be surprisingly stable under the vibrational perturbations typically studied in chemistry (with only $2$ and $3$ H-bonds identified as unstable for the Cage and Prism).
  • The saddles defining H-bond paths that are unstable under vibrations are also the saddles associated to the lowest persistence (Fig. 8). This indicates that low-persistence saddle pairs are likely to be canceled first by vibrations. Still in principle, path instability can also be observed for persistent saddles (see Fig. 4).
  • Our experiments (Fig. 6) suggest that the disappearance of an H-bond path under vibrations is gradual, involving first a mis-connection (that may be related to documented oxygen-oxygen contractions [84]), followed by a cancellation of its saddle.
  • Our experiments on documented proton tunneling (Figs. 9, 10) confirm the brevity of H-bond reconfiguration under this mechanism [65]. Moreover, this analysis identifies clear breaking points in terms of (Ac--H) bond length ($2.3$) and (Ac--Dn, Ac--H) bond angle ($14$°), beyond which a H-bond path disappears.
  • Finally, the geometrical analysis, at equilibrium, of the H-bond paths identified as unstable under vibrations (Fig. 13) reveals a correlation between our stability measure (bond occurrence rate, Sec. 4.4) and the above geometrical indicators. In particular, the most unstable H-bond paths are also the bonds whose indicators are the closest to the above limit values defining a breaking point. Then, the indicator ranges for unstable H-bond paths – bond length: ($2$, $2.3$) Å, bond angle: ($8$, $14$)° – can be interpreted as uncertainty intervals, within which H-bond paths are unstable and are likely to disappear under molecular vibrations (indicating modes that can contribute to isomerization of (H2O)6, Fig. 12).
We believe our work opens several research avenues at the intersection between DMT and QTAIM. First, our BondMatcher framework can be used for other studies requiring a robust, per-bond comparative analysis of several systems. For instance, the protocol used in our case studies could be extended to more complex molecular systems (e.g., as in [5]), to get a refined picture of the instability of H-bond paths. Second, our framework could also be used for the comparative analysis of a single molecular system, but studied under various descriptors (e.g., electron density versus bare nuclear potential). Such feature correspondences between descriptors could be instrumental to accelerate several high-throughput chemical analysis tasks, as the electron density can still be computationally expensive to obtain. However, such comparisons would require to revisit our extremum graph matching procedure, in particular to reliably account for distinct extremum counts.
Acknowledgements.
This work is partially supported by the European Commission grant ERC-2019-COG “TORI” (ref. 863464, https://erc-tori.github.io/) and the Polish National Science Centre (NCN) (grant 2020/38/E/ST4/00614). We gratefully acknowledge the Poland’s high-performance Infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no PLG/2024/017638. We also thank Prof. Dr. Jeremy Richardson (ETH Zürich, Switzerland) for providing molecular structures of the (H2O)6 prism isomer modeling the tunneling reactions.